OpenStudy (anonymous):

Suppose there are two cubes, with the first cube having a side of h, and the second cube having a side of 5h. What is the relation of the volume of the first cube to the volume of the second? A. V2 = 25V1 B. V2 = 125V1 C. V2 = 5V1 D. V2 = 5hV1

OpenStudy (anonymous):

In such questions compare the Volumes with the following f ormula: V =L^3 ( L = side of cube)

OpenStudy (anonymous):

V1 = H^3 and V2=(5h)^3 (V1\V2)=(h^3\5h^3) = 1:125

OpenStudy (anonymous):

For this, we should find volume of both cubes. Vol. of first cube = h * h * h = h^3 Vol. of second cube = 5h * 5h * 5h = 125h^3 So volume of second cube is 125 times greater than second one.

OpenStudy (anonymous):

sachinrajsharma has provided the correct solution.

OpenStudy (anonymous):

so its b?

OpenStudy (anonymous):

Yes its b.

OpenStudy (anonymous):

thank you.